### v2

# SciRIFF: A Resource to Enhance Language Model Instruction-Following over Scientific Literature

Abstract
We present SciRIFF (Scientific Resource for Instruction-Following and Finetuning), a dataset of 137K instruction-following demonstrations for 54 tasks covering five essential scientific literature understanding capabilities: information extraction, summarization, question answering, claim verification, and classification. SciRIFF demonstrations are notable for their long input contexts, detailed task specifications, and complex structured outputs. While instruction-following resources are available in specific domains such as clinical medicine and chemistry, SciRIFF is the first dataset focused on extracting and synthesizing information from research literature across a wide range of scientific fields. To demonstrate the utility of SciRIFF, we develop a sample-efficient strategy to adapt a general instruction-following model for science by performing additional finetuning on a mix of general-domain and SciRIFF demonstrations. In evaluations on nine held-out scientific tasks, our model—called SciTülu—improves over a strong LLM baseline by 28.1% and 6.5% at the 7B and 70B scales respectively, while maintaining general instruction-following performance within 2% of the baseline. We are optimistic that SciRIFF will facilitate the development and evaluation of LLMs to help researchers navigate the ever-growing body of scientific literature. We release our dataset, model checkpoints, and data processing and evaluation code to enable further research.
[Uncaptioned image] https://github.com/allenai/SciRIFF
[Uncaptioned image] https://huggingface.co/datasets/allenai/SciRIFF
1 Introduction
Large language models (LLMs) have the potential to advance scientific progress by helping researchers navigate and draw insights from the scientific literature. To accomplish these tasks, LLMs must be able to reliably follow a range of instructions—e.g. to extract information, summarize content, or answer questions—when given research articles as input. These instructions will often feature long input contexts, such as an entire research article. In addition, the model’s responses may need to be structured according to a specific format or schema that supports aggregation for literature review (Marshall and Wallace, 2019), or is consumable by software components like augmented reading interfaces (Lo et al., 2023; Palani et al., 2023). While bespoke models are available for specific scientific literature understanding tasks, models that can flexibly follow instructions are preferable both for their ease of use (offering a unified input / output interface) and for their ability to generalize to novel applications and settings.
The general instruction-following capabilities of LLMs have advanced rapidly in recent years, largely due to the availability of general-purpose instruction datasets (Zhang et al., 2023a). In addition, some instruction-following resources are available for specific scientific and medical tasks, such as describing the properties of a molecule (Fang et al., 2024; Yu et al., 2024) or answering medical exam questions (Toma et al., 2023; Han et al., 2023) (see §5 for a review). However, there is a scarcity of resources aimed at enabling flexible scientific literature understanding capabilities across a range of domains.
In this work, we present SciRIFF (Scientific Resource for Instruction-Following and Finetuning), a dataset to enable progress on instruction-following over scientific literature. SciRIFF includes 137K demonstrations for 54 tasks spanning five scientific literature understanding task categories: information extraction, summarization, question answering, claim verification, and classification. SciRIFF covers five scientific domains, ranging from artificial intelligence to clinical medicine (Figure 2). The tasks in SciRIFF are derived from existing scientific literature understanding datasets with human-annotated inputs and outputs, and are converted to a common instruction-following format via templates written by the paper authors (Figure 1). Many of the tasks feature long input contexts and require structured model responses.
We hold out nine representative tasks from SciRIFF for use as an evaluation benchmark, which we call SciRIFF-Eval3.1). We then perform supervised finetuning experiments to identify a sample-efficient strategy to adapt Tülu V2 (Ivison et al., 2023)—a strong open instruction-following model—for scientific literature use cases. We find that, by starting from the original Tülu V2 model and performing additional finetuning on a downsampled mix of SciRIFF and data from the Tülu V2 Mix, we are able to match the performance achieved by training on all instances, while using less than 20% of the available data. Using this sample-efficient training strategy, we improve performance on SciRIFF-Eval by 28.1% over a directly comparable baseline at 7B scale, and by 6.5% at 70B scale. At the same time, we achieve performance within 2% of the baseline model on a general instruction-following benchmark (§4.1). We publicly release our 7B and 70B models, which we call SciTülu.
In summary, our contributions are as follows:
  • We introduce SciRIFF, a dataset with 137K instruction-following demonstrations covering 54 literature understanding tasks spanning five scientific domains. Many tasks in SciRIFF feature long input contexts and require structured model responses.
  • We employ a sample-efficient approach to adapt a family of general instruction-following models to scientific literature use cases. The resulting SciTülu models achieve substantial performance gains on held-out scientific tasks, without sacrificing general capabilities.
  • We release the SciRIFF dataset, SciTülu model checkpoints, and code to recreate the dataset and perform evaluations on nine held-out tasks from SciRIFF.
Figure 1: Example SciRIFF tasks. Given an input context from a research paper, the  text prompt  instructs an LLM to perform an operation on the input—e.g. determine whether the  abstract  entails a scientific  claim , extract information over the  full_text , answer a question, etc. The model’s  output  must conform to a task-specific, user-specified  structure . SciRIFF unifies 54 scientific literature understanding tasks under a common input / output format, enabling the development of LLMs that can flexibly generalize to novel scientific use cases.
2 SciRIFF
SciRIFF is a comprehensive instruction-tuning resource focused on real-world scientific literature understanding, consisting of 137k high-quality instructions derived from 54 datasets, which span five task categories and five subject domains. The primary design objective of SciRIFF is to enhance and evaluate instruction-following capabilities of LLMs in this specialized domain. Our focus is on document-grounded scientific literature understanding tasks, rather than tasks that evaluate general reasoning or mathematical problem-solving abilities without reference to scientific literature (e.g., MMLU (Hendrycks et al., 2021a)). In addition to coverage of a wide range of tasks, the instructions in SciRIFF often are grounded in long input contexts (i.e., scientific papers), and they support structured outputs according to a specific schema useful for tasks in literature understanding (such as relation extraction, fact checking with rationale selection, question answering with attribution, etc). The instances in SciRIFF are sourced from existing high-quality scientific datasets and converted into instructions using human expert-written instruction templates. Below we discuss each of the steps involved in the creation of SciRIFF.
2.1 Dataset construction
We create SciRIFF by repurposing existing scientific literature understanding datasets for instruction-following—similar to resources like Flan (Longpre et al., 2023) and Super-NaturalInstructions (Wang et al., 2022)—rather than the alternative recent trend of generating synthetic data using an LLM (e.g., Köksal et al. (2023); Li et al. (2023)). We make this choice for two reasons. First, we believe it is sensible to exhaust available human-annotated resources, which we can be confident are correctly-annotated, before turning to potentially noisy synthetic data. Second, given the lack of existing evaluation benchmarks for scientific literature understanding, it may be difficult to assess the utility of synthetic data generation approaches. For this reason, we hold out nine SciRIFF tasks as an evaluation benchmark (§3.1), which we hope will provide valuable signal for future synthetic data generation efforts.
Dataset selection criteria
In forming SciRIFF, we focus on scientific literature understanding tasks in which the model is given a portion of scientific text as input, and is instructed to produce output derived directly from the text. This task family includes summarization, reading comprehension, information extraction and other tasks, and is the most relevant setting for real-world use cases (e.g., meta-analysis of literature, clinical decision-making, augmented reading). We exclude datasets that require retrieval from document collections (e.g., open-domain QA) and datasets that assess reasoning and mathematical problem-solving skills without necessarily relying on scientific literature, such as ScienceQA (Lu et al., 2022), SciBench (Wang et al., 2023a), and MATH (Hendrycks et al., 2021b). Additionally we only keep datasets that are publicly available and have a permissive license. Finally we include datasets that are well-documented and actively maintained. See Appendix A.1 for the complete task list.
Instruction templates
We convert our 54 datasets into natural language input-output pairs suitable for instruction tuning using dataset-specific instruction templates [^1] We used a single template per dataset. Future work could explore the utility of multiple templates. created by the paper authors, who are NLP experts. [^2] We conducted initial experiments using an LLM to create templates, but found that the resulting instructions were often vague and did not clearly specify the desired output format.
We use json as the common output format for all structured tasks, which facilitates consistent evaluation and matches with larger industry trends to request JSON model outputs. Instruction templates are written in Jinja (Pallets, 2024). Guidelines and best practices for prompt-writing are detailed in our GitHub repository. Each prompt was double-checked by an additional author for clarity and correctness.
2.2 Instruction Mix Statistics
(a) Task distribution
(b) Domain distribution
Figure 2: Distribution of task categories and domains in SciRIFF. The numbers in the pie charts indicate the number of datasets present in every task category/domain, while the numbers in brackets indicate the total number of instances per task category/domain.
Figures 2(a) and 2(b) present an overview of the SciRIFF training set distribution over task categories and domains respectively. Given the significant presence of information extraction tasks, it is unsurprising that a large percentage of datasets in SciRIFF (34 datasets; 63%) require structured outputs.
We construct three instruction mixes from this dataset collection, with maximum context lengths (input + output tokens) of 4,096, 8,192 and 16,382 per instance (longer instances are truncated where possible and discarded otherwise; see Appendix A.3). Due to model and hardware limitations, we conduct experiments in this work using the SciRIFF-4096 mixture, and make the longer mixtures available to enable future research. In what follows, we refer to SciRIFF-4096 simply as SciRIFF.
Input prompts in SciRIFF average over 1,200 tokens in length (Figure 5). As shown in Appendix A.2, this is over three times longer than the average prompt length for a range of representative general-purpose and domain-specific instruction datasets; for instance, Flan prompts average 350 tokens.
3 Experimental setup
We conduct finetuning experiments to determine the best way to use SciRIFF to improve LLM performance on scientific instruction-following.
3.1 Evaluation
Science evaluation
We hold out 9 tasks from SciRIFF for evaluation, covering a representative range of task categories and scientific domains. The inputs, outputs, and evaluation metrics for each task are detailed in Table 1. Additional details along with full input / output examples for all evaluation tasks are included in Appendix D.
Table 1: Evaluation tasks included in SciRIFF-Eval. “/” separators indicate two separate subtasks. We use GPT-3.5 Turbo as our LLM judge and evaluate similarity on a 1-5 scale; see Appendix D.
Name Input Output Metrics
BioASQ List QA Question over a set of paper excerpts List of answer entities Exact match F1
BioRED Biomedical research abstract Mentions of 6 biomedical entity types Exact match F1
DiSCoMaT LaTex excerpt containing a table Values of table entries BLEU score vs. reference values
Evidence Inference (EI) Clinical trial report abstract (intervention, comparator, outcome, effect, evidence) tuples String overlap approximate match F1
MultiCite (MC) Citation context List of expressed citation intents (7 categories) Exact match F1
MuP Machine learning paper full text Short summary useful for a peer review Similarity to reference summary, evaluated by LLM judge
Qasper Question over an NLP paper full text Answer to question / Paragraphs providing attribution Similarity to reference answer, evaluated by LLM judge / Token F1 vs. reference paragraphs
SciERC Computer science abstract Mentions of 6 computer science entity types Exact match F1
SciFact Claim to verify against a biomedical research abstract Fact-checking verdict / Evidentiary sentences Label F1 / Token F1 vs. reference evidence
General instruction-following evaluation
We re-use the instruction-following evaluations from Ivison et al. (2023). These test a broad range of general abilities including world knowledge (MMLU), reasoning (GSM8k, Big Bench Hard), coding (CodexEval), open-ended generation (AlpacaEval), and truthfulness (TruthfulQA) [^3] We consider the toxicity and multilinguality evaluations from Ivison et al. (2023) as out-of-scope for this work and do not evaluate on them; hence our reported averages are not directly comparable Ivison et al. (2023)..
3.2 Finetuning approach
We experiment with two training data sources and two starting model checkpoints. In §4.2, we present ablation results examining the effects of both components on model performance.
Data sources
We finetune on combinations of two datasets. (1) SciRIFF. We train on a fixed number of instances per task $n_{sci}$ by sampling from each task at random; we use $n_{sci}=1000$ for our final models based on the results of ablation experiments (§4.2). For tasks with fewer than $n_{sci}$ instances, we include all examples. (2) Tülu V2 Mix (Ivison et al., 2023). Tülu V2 Mix  [^4] We generally use “Tülu V2” to refer to the model and “Tülu V2 Mix” to refer to the dataset. In cases where it is clear from context, we may refer to the dataset simply as “Tülu V2”. is a high-quality general instruction-following dataset. It includes demonstrations from a range of sources, both human-written (e.g. Flan (Wei et al., 2022)) and distilled from proprietary LLMs (e.g. ShareGPT [^5] https://sharegpt.com/, Open Assistant [^6] https://github.com/LAION-AI/Open-Assistant).
Starting checkpoints
We conduct finetuning experiments starting from two model checkpoints: (1) Llama 2. We train on all available Tülu V2 Mix demonstrations, combined with $n_{sci}$ instances per SciRIFF task. (2) Tülu V2. We start from a Tülu V2 checkpoint already trained on the full Tülu V2 Mix, and perform additional finetuning on $n_{sci}$ instances per SciRIFF task, together with a matching number of instances randomly sampled from Tülu V2 Mix. Including Tülu V2 Mix instances prevents the model from “forgetting” its general instruction-following capabilities; see §4.2. Table 2 shows the number of training examples used to finetune each starting checkpoint.
Table 2: SciRIFF and Tülu V2 Mix  [^7] The original Tülu V2 Mix contains 326,154 examples, including 7.5K scientific literature understanding demonstrations. We remove these for our experiments. instances used for finetuning, with $n_{sci}=1000$.
Checkpoint SciRIFF Tülu-V2 Total
Llama 2 base 35,357 318,686 354,043
Tülu V2 35,357 35,357 70,714
SciTülu
For our final 7B and 70B models, we use Tülu V2 as our starting checkpoint and set $n_{sci}=1000$. This approach leads to performance comparable to finetuning Llama base, while requiring 80% less data (Table 2). We call our final models SciTülu.
4 Results
We present the main results for SciTülu in §4.1. In §4.2, we perform ablations examining the effects of the different finetuning data sources and model checkpoints.
4.1 Main Results
Results for our final model, SciTülu, are shown in Table 3. Tülu V2 serves as a directly comparable baseline. We also include the performance of widely used open (Llama 2 7B Chat and Llama 8B Instruct) and closed (GPT-3.5 and GPT-4888Predictions were made using gpt-3.5-turbo-1106 and gpt-4-turbo-0125-preview.) LLMs.
SciRIFF greatly improves scientific performance at 7B, with moderate improvement at 70B.
SciTülu achieves a 28.1% average improvement on SciRIFF-Eval over Tülu V2, while achieving nearly identical performance on the general evaluations. Llama 3 exhibits the strongest performance on the general evaluations but underperforms SciTülu on SciRIFF-Eval by 20%. Llama 2 has the lowest performance on both evaluations. [^9] During manual inspection, we found that Llama 2 often struggles to generate properly-structured responses, leading to very low scores on some tasks. Future work could leverage tools like jsonformer to alleviate this issue. At the 70B scale, training on SciRIFF provides a 6.5% average improvement on SciRIFF-Eval while causing a slight 2.1% decrease on general tasks. The smaller SciRIFF-Eval performance gain may be due to the stronger reasoning capabilities of 70B models, which can more readily perform complex tasks zero-shot.
SciTülu 7B matches Tülu V2 70B on science.
Given that many scientific literature understanding workflows (e.g. literature review) require running an LLM over large research corpora, the ability to run inference using a 7B model in place of a 70B could provide substantial efficiency gains and cost savings to practitioners.
Table 3: Performance of SciTülu. Tülu V2 serves as a directly comparable baseline. Popular open (Llama) and closed (GPT) instruction-following models are also shown for reference. We report performance on each task in SciRIFF-Eval, along with the average performance on our general-purpose evaluations. The last two columns are averages. Complete results for general-purpose metrics can be found in Appendix C. Columns with a “/” indicate two evaluation metrics as described in §3.1. By training on SciRIFF, SciTülu improves over Tülu V2 by 28.1% at 7B, and 6.5% at 70B.
Model BioASQ BioR DiscMT EI MC MuP     Qasper SciERC    SciFact Sci. Gen.
GPT-3.5T 47.3 53.9 67.9 19.2 47.8 76.8 54.7 / 39.8 28.6 69.7 / 53.3 50.8 76.2
GPT-4 46.7 61.0 78.3 24.7 58.7 86.9 67.8 / 50.5 42.2 84.3 / 68.7 60.9 88.8
Llama 2 7B chat 34.2 0.0 4.8 7.4 37.8 72.0 15.7 / 8.5 0.3 27.7 / 6.2 19.5 36.4
Llama 3 8B instruct 43.3 40.3 37.3 13.5 37.9 84.6 41.1 / 25.9 25.4 42.3 / 40.1 39.2 72.9
Tülu V2 7B 44.5 15.1 47.8 15.1 33.6 71.6 43.2 / 26.3 21.2 49.5 / 35.4 36.7 47.8
SciTülu 7B (ours) 37.5 55.7 61.5 11.6 34.6 72.1 54.2 / 38.6 35.6 66.0 / 49.2 47.0 47.5
Tülu V2 70B 38.3 50.6 68.2 16.9 48.5 64.9 49.1 / 20.7 32.5 76.4 / 57.2 47.6 69.8
SciTülu 70B (ours) 42.7 69.3 72.6 17.5 62.8 61.1 43.0 / 27.3 35.9 70.7 / 55.6 50.7 68.3
Gains at 7B are largest on tasks requiring extraction and attribution.
SciTülu 7B outperforms Tülu V2 most substantially on information extraction tasks (BioRED, DiSCoMaT, and SciERC) [^10] SciTülu does not improve on Evidence Inference, which is an IE task. This likely occurs because the task is very difficult (Appendix D) and evaluation is challenging (§6), making current evaluations unreliable. Improving modeling and evaluation for complex IE tasks represents important future work. and tasks requiring attribution (Qasper, SciFact), achieving a 51% average improvement. On the other hand, it performs worse on summarization (MuP) and question answering (BioASQ). This may occur because summarization and QA are well-covered by existing instruction-following resources, while tasks like information extraction and attribution have received less attention.
GPT-4 achieves the strongest performance, but has room to improve.
GPT-3.5 Turbo performs roughly on par with SciTülu on SciRIFF-Eval. Interestingly, GPT-4 performance is still relatively low on many tasks; this is due to a combination of task difficulty and evaluation challenges, which we discuss in §6.
4.2 Ablations
Table 4: Ablations examining the effect of starting checkpoint and training data on scientific and general instruction-following. For Data, SciRIFF and Tülu each train on one of the two data sources described in §3.2, while SciRIFF + Tülu trains on both. All settings use $n_{sci}=1000$. See Appendix C for individual task metrics.
Starting Checkpoint Finetuning data Sci Avg. Gen. Avg. Sci Avg. Gen. Avg.
Model size=7B Model size=70B
Llama 2 base Tülu 36.7 47.8 47.5 69.8
SciRIFF 48.0 23.9 51.1 44.2
SciRIFF+Tülu 46.0 48.9 50.8 70.4
Tülu V2 SciRIFF 47.0 33.3 48.8 56.9
SciRIFF+Tülu 47.0 47.5 50.7 68.3
Figure 3: Performance on SciRIFF-Eval as a function of $n_{sci}$, the number of science instances per training task. Performance gains largely saturate by $n_{sci}=1000$. Experiments are conducted on 7B models.
We conduct ablations to characterize the effects of (1) the inclusion of each data source, (2) the choice of starting checkpoint, and (3) the number of instances per SciRIFF task.
Table 4 reports average SciRIFF-Eval and general metrics for our two starting checkpoints using three data configurations: (1) Tülu trains on the Tülu V2 mix with no science instructions, (2) SciRIFF trains only on science instructions, and (3) SciRIFF + Tülu represents our final training mix described in §3.2.
Training on SciRIFF+Tülu gives the strongest overall performance
Models trained only on SciRIFF perform well on science evaluations but struggle at general instruction-following. The Llama-2 checkpoint trained only on Tülu performs well on general tasks but is weaker on science. Training on the combined mix achieves the best overall performance on our general metrics and achieves comparable science performance to training only on SciRIFF, with slight increases in some settings (e.g. 70B Tülu V2) and slight decreases in others (e.g. 7B Llama 2 base).
Starting from Tülu V2 performs comparably to Llama 2 base, while using 20% of the compute
When trained on SciRIFF+Tülu data, models trained starting from Tülu V2 vs Llama 2 base are comparable. Tülu V2 is slightly better on science at 7B and nearly identical at 70B, while it underperforms Llama 2 by less than 3% on general evals. Given that finetuning Tülu V2 requires only 20% of the data (§3.2), we recommend that practitioners save on compute by instruction-tuning a strong instruction-following model on a mix of SciRIFF and general-domain instructions. This aligns with findings from prior works, e.g. Dong et al. (2024); Shi et al. (2023).
1,000 instances per science task is sufficient for peak performance
Figure 3 shows that performance on SciRIFF-Eval increases sharply as $n_{sci}$ rises from 100 to 500 and levels off subsequently. We set $n_{sci}=1000$ based on these findings.
5 Related Work
Strategies for creation of instruction-following resources.
Instruction tuning, or finetuning LLMs to improve their instruction-following ability, has emerged as a crucial technique for enhancing generalizability and controllability of LLMs  (Wei et al., 2022; Sanh et al., 2022; Mishra et al., 2022; Ivison et al., 2023). Several strategies have been explored for the creation of instruction-following resources, such as repurposing existing datasets using human-written instruction templates (Wei et al., 2022; Chung et al., 2024; Sanh et al., 2022; Mishra et al., 2022), crowdsourcing instructions [ Databricks (2023); Zhou et al. (2023), ShareGPT [^11] https://sharegpt.com/] and using LLMs to generate synthetic instructions data. As LLM capabilities grow stronger, synthetic instruction generation approaches, often including humans in the loop as correctors, have shown promising results. Broadly, these approaches use LLMs to either generate new dataset/task instances alongside instructions (Wang et al., 2023c; Xu et al., 2024; Nayak et al., 2024; Lou et al., 2024), or to “back-translate” existing datasets into instructions (Yin et al., 2023; Köksal et al., 2023; Li et al., 2023). In this work, we create instructions using human-written templates (§2.1).
Instruction-following resources for scientific literature.
While numerous open-domain instruction-following collections exist, resources for enhancing and evaluating LLMs’ instruction-following capabilities on scientific literature are limited. Such resources are critical for developing models that can assist researchers and accelerate scientific discovery (Taylor et al., 2022; Xie et al., 2023). Recent work has taken steps in this direction with the development of instruction-following datasets for specific domains such as mathematics (Yue et al., 2024a, b; Shao et al., 2024; Luo et al., 2023; Tang et al., 2024; Toshniwal et al., 2024), medicine (Parmar et al., 2022; Wu et al., 2024; Rohanian et al., 2023), chemistry (Yu et al., 2024; Zhang et al., 2024b), molecular biology (Fang et al., 2024; Tran et al., 2023), materials science (Song et al., 2023), and college-level foundational science  (Zhang et al., 2024a). Besides domain limitations, these resources primarily focus on improving LLMs’ abilities to solve college-level science problems or reasoning tasks (see also, MMLU (Hendrycks et al., 2021a), SciEval (Sun et al., 2023), TheoremQA (Chen et al., 2023), SciBench (Wang et al., 2023a), and GPQA (Rein et al., 2023)). In contrast, SciRIFF both covers a broader set of scientific domains and focuses on document-grounded scientific literature understanding tasks that can power real-world scientific use cases. Another distinguishing factor of our work is our inclusion of tasks that require structured outputs, following a uniform JSON output format, besides text-to-text tasks. Some instruction-tuning resources have explored structured output formats  (Zhang et al., 2023b; Wang et al., 2023b; Jiao et al., 2023; Gao et al., 2023), but not with a focus on scientific literature. Finally, most datasets in SciRIFF require long input contexts, leading to longer instruction contexts than prior work (see Appendix Table 5 for a comparison).
Other scientific literature benchmarks.
In addition to instruction-following resources, prior works have also developed benchmarks to improve and assess scientific literature understanding. Notable efforts in the biomedical domain include BLUE (Peng et al., 2019), BLURB (Gu et al., 2021), InBoXBART (Parmar et al., 2022), and BigBio (Fries et al., 2022); SciRIFF covers a broader set of domains than these resources. Other efforts such as SciRepEval (Singh et al., 2023), Galactica (Taylor et al., 2022), and AcademicGPT (Wei et al., 2023) cover domains beyond biomedicine, but are not suitably formatted for training or evaluating instruction-following models.
6 Conclusion and Future Work
In this work, we introduced SciRIFF, a resource to facilitate progress on LLM instruction-following over scientific literature. We demonstrated that training on SciRIFF leads to improved performance on held-out scientific tasks, with especially large improvements at the 7B scale on tasks requiring structured extraction or attribution.
Our work points toward a number of future research directions. As observed in §4.1, GPT-4 performance on SciRIFF-Eval is fairly low. This is partly due to the difficulty of the tasks, but also due the challenges associated with evaluating structured LLM responses in cases where the predicted surface form does not match the reference, but the underlying meaning is the same (Wadhwa et al., 2023). Utilizing LLMs to perform more flexible evaluations (Kim et al., 2024) represents a promising direction. Future work could also explore whether diversifying tasks by using multiple templates—potentially with different input and output formats—could enable models to learn from a larger number of demonstrations and to generalize more readily to unseen tasks. Synthetic data generation techniques may also be beneficial.
In conclusion, we are optimistic that the SciRIFF data and evaluations, as well as the SciTülu models, will serve as valuable resources to build systems which can boost the productivity and creativity of scientific researchers.
7 Acknowledgements
We thank Charlie Gao, Minyi Chen, Yujie Qiao, and Kaiyuan Guan who helped with the initial stages of our dataset creation. We are grateful for the compute support provided by the Google TRC program.
Appendix A Additional information on SciRIFF
A.1 SciRIFF task list
The full list of SciRIFF tasks is visualized in Figure 4. Detailed information on all tasks—including citations, URLs to source websites, and licensing information where available—is provided in our dataset card, available at https://huggingface.co/datasets/allenai/SciRIFF. Where convenient, we use datasets as preprocessed by the BigBio resource (https://huggingface.co/bigbio); details are in the dataset card.
[…]
A.2 Task length distribution
Figure 5: Distribution of input (left) and output (right) token lengths over SciRIFF training instances.
Figure 5 shows the distribution of input and output lengths for demonstrations in SciRIFF.
Table 5: Comparison with selected instruction-following datasets. We use the following abbreviations: PH – Physics; FP – Formal Proof; MatSci – Materials Science. We report average token counts for input/output using Llama 2 tokenizer using up to 200k subsamples from each dataset. BoX dataset is not readily available.
Name # Instances Domain Avg. Length
General Domain
Flan V2 [Chung et al., 2024] 15M General 355.6 / 31.2
SuperNI [Wang et al., 2022] 97K General 291.1 / 38.7
Tülu V2 Mix [Ivison et al., 2023] 326K General 353.3 / 696.9
Scientific Domain
BoX [Parmar et al., 2022] 141K Biomed X
SciInstruct [Zhang et al., 2024a] 254K Math, PH, Chem, FP 88.4 / 265.6
Mol-Instructions [Fang et al., 2024] 2.04M Biomolecular 126.3 / 112.9
MathInstruct [Yue et al., 2024a] 262K Math 82.5 / 174.0
MedInstruct-52K [Zhang et al., 2023c] 52K Medical 148.2 / 96.9
LlaSMol [Yu et al., 2024] 3.29M Chem 81.9 / 53.0
SciRIFF (Our work) 137K AI, Biomed, Clinical, Chem, MatSci 1242.9 / 139.6
Table 5 compares SciRIFF with selected instruction-following datasets, including canonical collections commonly used for general fine-tuning and selected recent datasets specialized in scientific domains. Our dataset features longer input contexts than existing resources.
A.3 Truncation strategy
In §2.2, we mention that when an instance exceeds the maximum context length for a given version of SciRIFF, we truncate where possible and discard otherwise. In particular, we truncate for tasks (like question answering) where the task output can be localized to particular passages in the input document by randomly removing irrelevant passages until the document fits in the desired context. For tasks like summarization, where the task output cannot easily be localized, we simply discard examples that are longer than the context window.
Appendix B Training Details
For instruction-tuning, our training hyperparameters were as follows:
  • Precision: BFloat16
  • Epochs: 2
  • Weight decay: 0
  • Warmup ratio: 0.03
  • Learning rate: 2e-5 (1e-5 for 70B)
  • Max. seq. length: 4,096
  • Effective batch size: 128
All of our models were trained on v3-128 TPUs on the Google TPU Research Cloud.
Appendix C Full modeling results
Table 4 in §4.2 reports average science and general evaluation metrics for our ablations. Here, we include individual task metrics. Science metrics are shown in Table 6. General evaluation metrics are shown in Table 7.
Table 6: Model ablation results on science tasks.
Checkpoint Data BioASQ BioRED DiSCoMaT Ev. Inf. MultiCite MUP QASPER SciERC SciFact
7B
Llama 2 Tülu 44.5 15.1 47.8 15.1 33.6 71.6 43.2 / 26.3 21.2 49.5 / 35.4
Llama 2 SciRIFF 36.1 61.0 62.0 4.2 54.4 70.9 49.8 / 45.2 35.6 60.9 / 47.6
Llama 2 SciRIFF+Tülu 27.0 50.1 60.6 11.8 44.1 70.9 50.0 / 44.4 34.5 65.2 / 47.4
Tülu V2 SciRIFF 39.1 55.6 64.5 0.0 52.2 67.8 48.6 / 38.2 35.7 67.3 / 47.5
Tülu V2 SciRIFF+Tülu 37.5 55.7 61.5 11.6 34.6 72.1 54.2 / 38.6 35.6 66.0 / 49.2
70B
Llama 2 Tülu 38.3 50.6 68.2 16.9 48.5 64.9 49.1 / 20.7 32.5 76.4 / 57.2
Llama 2 SciRIFF 39.2 68.1 71.6 14.2 64.0 62.5 43.9 / 27.1 43.8 73.3 / 54.5
Llama 2 SciRIFF+Tülu 36.5 59.7 72.8 18.6 63.7 64.9 45.1 / 26.6 37.0 77.5 / 56.9
Tülu V2 SciRIFF 12.5 66.8 69.1 14.4 56.8 70.7 41.1 / 25.0 41.0 81.4 / 58.0
Tülu V2 SciRIFF+Tülu 42.7 69.3 72.6 17.5 62.8 61.1 43.0 / 27.3 35.9 70.7 / 55.6
Table 7: Model ablation results on general tasks.
BBH Codex-eval GSM8K MMLU TruthfulQA AlpacaEval Average
Checkpoint Data 3-shot CoT Pass@10 8-shot CoT 0-shot
Info
* True
% Win vs
Davinci-003
7B
Llama 2 Tülu 45.6 37.7 29.5 50.2 50.4 73.1 47.8
Llama 2 SciRIFF 38.4 18.7 12.0 39.5 27.8 6.7 23.9
Llama 2 SciRIFF+Tülu 43.8 36.6 34.0 49.7 54.3 74.7 48.9
Tülu V2 SciRIFF 42.5 29.0 23.0 45.8 41.0 18.8 33.3
Tülu V2 SciRIFF+Tülu 42.7 35.9 29.5 49.6 52.6 74.7 47.5
70B
Llama 2 Tülu 66.8 67.6 66.5 67.7 63.8 86.6 69.8
Llama 2 SciRIFF 54.4 44.5 53.5 65.3 36.1 11.2 44.2
Llama 2 SciRIFF+Tülu 68.1 67.3 73.5 66.8 60.7 86.1 70.4
Tülu V2 SciRIFF 66.2 57.2 67.0 66.3 49.9 34.5 56.9
Tülu V2 SciRIFF+Tülu 67.0 60.3 67.5 66.8 62.2 85.8 68.3
Appendix D Evaluation details
The following pages show full input / output examples for all SciRIFF-Eval tasks, along with details on metric calculations. This information is also available on the project GitHub page.
See pages - of docs/evaluation_tasks.pdf